ANALYSIS OF HIGH-RESOLUTION INFRARED SPECTRA OF ^{11}BF_{3} FOR VIBRATIONAL STATES BE-TWEEN 1600 AND 4300 cm^{−1}

ARTHUR MAKI, 15012 24th Ave. S.E., Mill Creek WA, 98012; TONY MASIELLO, \underline{THOMAS A. BLAKE}, Pacific Northwest National Laboratory, P.O. Box 999, Mail Stop K8-88, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC05-76RL0 1830).

Last year at this Symposium (RX02) we presented spectroscopic measurements and ro-vibrational analysis for vibrational states of ^{11}BF_{3} below 1600 cm^{−1}. This year we present measurements and analysis for vibrational states of ^{11}BF_{3} up to 4300 cm^{−1}. Measurements were made of an isotopically enriched sample using a Bruker IFS 120HR Fourier transform spectrometer located at the Pacific Northwest National Laboratory. Spectra were recorded with resolutions ranging from 0.0015 to 0.0035 cm^{−1} and pathlengths up to 32 m. The combination states in the following sets of interacting states have been either observed directly from a transition or determined indirectly by its perturbative affects on observed states: 102^{2}0^{0} (3783.85162(8) cm^{−1}), 102^{0}0^{0} (3756.085 cm^{−1}), 101^{−1}3^{3} (3763.14(16) cm^{−1}), 002^{0}2^{2} (3830.233(50) cm^{−1}); 200^{0}1^{1} (2240.94976(3) cm^{−1}), 120^{0}0^{0} (2264.327(10) cm^{−1}); 300^{0}1^{1} (3118.20602(6) cm^{−1}), 220^{0}0^{0} (3141.688(13) cm^{−1}); 110^{0}1^{1} (2050.11053(7) cm^{−1}), 030^{0}0^{0} (2081.12683(6) cm^{−1}); 010^{0}2^{0} (1652.35840(7) cm^{−1}), 010^{0}1^{2} (1652.73764(5) cm^{−1}); 101^{1}0^{0} (2336.2009(29) cm^{−1}), 100^{0}3^{1,3} (2311.519(15) cm^{−1}); 201^{1}0^{0} (3216.2986(13) cm^{−1}), 200^{0}3^{1,3} (3188.650(76) cm^{−1}); (001^{1}1^{1})^{0,2} (1931.87377(14) cm^{−1}), (000^{0}4^{0,2,4}) (1921.996(11) cm^{−1}); (101^{1}1^{1})^{0,2} (2810.69018(8) cm^{−1}), (100^{0}4^{0,2,4}) (2787.31 cm^{−1}); (201^{1}1^{1})^{0,2} (3687.1503(10) cm^{−1}). The 003^{1}0^{0} − 000^{0}0^{0} transition was observed near 4310 cm^{−1} and was treated as an unperturbed perpendicular band except for the l-type resonance between the k = 1, l = 1 and k = −1, l = −1 levels. Effects from other perturbations are thought to be too small to be observed. An infrared forbidden transition 011^{1}0^{0} − 000^{0}0^{0}, E′′ − A′_{1}, was also observed near 2140 cm^{−1}. The transitions obey electric dipole allowed selection rules Δk = ±2, Δl = ∓1. Intensity comes from a term in the dipole moment operator that governs the intensity of the 002^{2}0^{0} − 000^{0}0^{0} transition and to a lesser extent from terms that govern the intensity of the 001^{1}1^{1} − 000^{0}0^{0} transition and the fundamentals. A detailed discussion of the interactions and the fit spectroscopic constants will be presented for the vibrational states as well as the Hamiltonian used to derive them.